OpenStudy (anonymous):

Find the arc length of the graph, on the interval [1/2,2], of y=((x^3)/(6))+((1)/(2x))

OpenStudy (anonymous):

Hi. the arc length is generally : L=integral[(1+y')^(1/2)]

OpenStudy (anonymous):

sorry i could not use symboles

OpenStudy (anonymous):

i still need help

OpenStudy (anonymous):

1+y'=1+(1/2)x^2+ (1/2) lnx

OpenStudy (anonymous):

and we should integral in our interval

OpenStudy (zarkon):

\[\int\limits_{a}^{b}\sqrt{1+\left(\frac{dy}{dx}\right)^2}dx\]

OpenStudy (anonymous):

i mean 33/16

OpenStudy (anonymous):

Um, shomething not right there...

OpenStudy (anonymous):

can you help me with this problem?

OpenStudy (zarkon):

\[\int\limits_{1/2}^{2}\sqrt{1+\left(\frac{d(x^3/6+1/(2x))}{dx}\right)^2}dx\] \[=\int\limits_{1/2}^{2}\frac{x^4+1}{2x^2}dx\]

OpenStudy (zarkon):

\[=\frac{33}{16}\]

OpenStudy (anonymous):

thanks